For the remainder of questions, assume that the sl
Answer #1

5) C

W = F.s = -umgCos theta1 *L1    negative because friction opposite to the direction of motion,\

F=uN=mgCos theta1

6)W = F.s

W= -umg Cos theta2 *L2

negative because friction opposite to the direction of motion,F=uN=umgCos theta2,

as N is normal raction =mgCos theta2

7)B,

by energy conservation

Initial Potential Energy = final potential energy + kinetic energy gained + energy lost due to friction

mgh1 = mgh2 + 0.5mv2 + umg Cos theta2 *L2   + umg Cos theta2 *L2  

mgh1 = mgh2 + 0.5mv2 - (- umg Cos theta2 *L2   - umg Cos theta2 *L2 )

mgh1 = mgh2 + 0.5mv2 - Wfriction    as work done by friction is negative